Solve this system using the Addition Method, describe in words the steps you will take to do that.
3x - y = -1
4x + 2y = 22

Answer by Alan3354(69443)

Solve this system using the Addition Method, describe in words the steps you will take to do that.
3x - y = -1
4x + 2y = 22
---------------
Multiply one or both eqns by coefficients to get the same coeff for one of the variables.
In this case, multiply the 2nd eqn by 1/2 to get 1y. Then add them to eliminate y, solve for x, etc.
3x - y = -1
4x + 2y = 22
--------------
3x - y = -1
2x + y = 11
------------ Add them
5x = 10
x = 2
-------------
Sub for x into either eqn to find y.
3x - y = -1
6 - y = -1
y = 7